OpenStudy (anonymous):

Solve for a: v^2 = v0^2 + 2a(delta)x

OpenStudy (anonymous):

Solve for a: v^2 = v0^2 + 2a(delta)x v^2 - v0^2 = 2a(delta)x divide both sides by 2(delta)x to get a....:D